# Beyond Full Label: Single-Point Prompt for Infrared Small Target Label Generation

### v1

@@ Introduction
Figure 4 illustrates that the results of EDGSP are closer to the target silhouette in the infrared image than those of other models. In fact, many pseudo masks such as Misc_28, Misc_74, 000595, XDU104, and XDU697 even exceed the quality of ground truth (GT). We also observe that COM and MCLC tend to merge the target with the background, while LESPS, MCGC, and Robust SAM struggle to accurately infer the morphology of infrared small targets. Additionally, SAM is incapable of labeling infrared small targets.
Figure 5: Quantitative results of three baselines at six cropping scales in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{at}}(\%)$. Dashed and solid lines represent label generation results from direct center-cropping and from using EDGSP on the cropped image, respectively.
@@
Table 3: Based on SCTransNet (SCT), ablation study of the single prompt embedding (SPE), TEI, rough edge map (REM), DPE, and BBM on SIRST, NUDT-SIRST, and IRSTD-1k. SPE means directly concatenating the single-point prompt with the original image as input. More details of each module are discussed in the supplementary material.
@@
Figure 5 showcases the results of three baselines (U-Net, DNA-Net, and SCTransNet) under prompt-based center-cropping. With the crop size increasing, model performance initially improves and then declines, and fails to achieve high-performance labeling (${P}_{d}<100\%$, IoU$<$85). This is because small crop sizes fail to encompass large targets, while large crop sizes introduce excessive irrelevant padding at the image edges. Additionally, target overlap is challenging to manage when splicing the inferred image back into the original image. In summary, missing boundary coordinate information diminishes the effectiveness of the formal center-cropping strategy. Therefore, EDGSP directly processes the entire image.
Table 4: Quantitative results of five IRSTD trained with five types of pseudo-labels and full labels on downstream test sets in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, ${{F}_{a}}(\%)$. The best and the second-best results are highlighted in bold and underlined, respectively.
@@
As shown in Table 3, SCTransNet is used as the baseline to validate the effectiveness of each proposed strategy. In M2, just incorporating single prompt embedding, the value of ${P}_{d}$ and IoU exceeds the full cropping strategies in Figure 5, illustrating the necessity of entire image processing. Using SPE, the IoU value further improved, and ${{P}_{d}}$ reached 100%, which demonstrated the SPE is capable of fleshing out target shapes and highlighting edge differences. The incorporation of the rough edge map raised the IOU further but led to target sticking. In M5, SPE is replaced with DPE. We can see that prompt embedding at the tail of the model enhances target edge differences again and improves overall performance. Finally, BBM accurately eliminates false alarms and EGDSP achieves accurate annotation. In Figure 6, with the help of target energy initialization, our model quickly responds to the target area and accurately predicts the left rotor of the drone. In scene two, without double prompt embedding (Ours w/o DPE), neighboring targets blend. We further observe that the attention map also responded in the background region, revealing that the local shape prediction remains connected to the global context modeling. This finding underscores again the importance of preserving the entire scene.
@@
We used the inferred pseudo masks [^4] The number of data is 1505, training and test sets split as 8:2, we performed three cross-validations. from the test sets of SIRST, NUDT-SIRST, and IRSTD-1k to train the detection model, to validate the effectiveness of the pseudo labels for downstream IRSTD. Quantitative results are shown in Table 4. EDGSP outperforms other single-point prompt label generation methods. Although MCGC achieves the lowest ${{F}_{a}}$, it also has a low IoU score. Remarkably, EDGSP not only reduces time cost by 80% but also outperforms full labels in detection task. Even with coarse point prompts, It still achieves 99.5% of the performance of full labels. Visual results are shown in Figure 7, compared to full labels, DNA-Net trained with EDGSP’s pseudo mask demonstrates enhanced shape inference and lower false alarm. This reaffirms the applicability of the proposed methodology.
Figure 7: Visual results of DNA-Net. “Point Result” and “Full Result” indicate the outcomes of DNA-Net trained with the pseudo label from EDGSP and the full label, respectively. The correctly detected targets and false alarms are highlighted in orange and yellow circles, while incomplete shape inferences are marked with arrows.
@@ Conclusion
In this article, we propose a novel annotation paradigm: enhancing detection networks with single-point prompts to accomplish the high-level IRSTLG. Specifically, we introduce the energy double guided single-point prompt (EDGSP) framework. It includes target energy initialization for effective shape evolution, double prompt embedding to prevent label adhesion, and bounding box-based matching to eliminate false alarms. This framework achieves accurate annotation on the SIRST, NUDT-SIRST, and IRSTD-1k datasets for the first time, and demonstrates superior performance compared to full labels in downstream IRSTD. In the future, we will focus on the coarse single-point prompt IRSTLG to further improve the utility of this framework.

### v4

@@ Introduction
Figure 4 illustrates that, compared to other models, the pseudo masks generated by EDGSP are closer to the target shape. In fact, many masks such as Misc_28, Misc_74, 000595, XDU104, and XDU697 even exceed the quality of ground truth (GT). We also observe that COM and MCLC tend to merge the target with the background, while LESPS, MCGC, and Robust SAM struggle to accurately infer the morphology of infrared small targets. Additionally, SAM is unsuitable for annotating IR small targets.
Figure 5: Quantitative results of three baselines at six cropping scales in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, and ${{F}_{at}}(\%)$. Dashed and solid lines represent label generation results from direct center-cropping and from using EDGSP on the cropped image, respectively.
@@
Table 3: Based on SCTransNet (SCT), ablation study of the single prompt embedding (SPE), TEI, rough edge map (REM), DPE, and BBM on mixed datasets from SIRST, NUDT-SIRST, and IRSTD-1k. SPE means concatenating the single-point prompt with the original image as input.
@@
Figure 5 showcases the results of three baselines (U-Net, DNA-Net, and SCTransNet) under prompt-based center-cropping. With the crop size increasing, model performance initially improves and then declines, and fails to achieve high-performance labeling (${P}_{d}<100\%$, IoU$<$85). This is because IR targets are multi-scale; small crop sizes fail to encompass large targets, while large crop sizes introduce excessive irrelevant padding at the image edges. Additionally, target overlap is challenging to manage when splicing the inferred image back into the original image. In summary, missing boundary coordinate information diminishes the effectiveness of the formal center-cropping strategy. Therefore, our EDGSP directly processes the entire image.
Table 4: Quantitative results of five IRSTD trained with five types of pseudo-labels and full labels on downstream test sets in terms of $IoU(\%)$, ${{P}_{d}}(\%)$, and ${{F}_{a}}(\%)$. The best and the second-best results are highlighted in bold and underlined, respectively.
@@
As shown in Table 3, SCTransNet is used as the baseline to validate the effectiveness of each proposed strategy. In M2, just incorporating single prompt embedding, the value of ${P}_{d}$ and IoU exceeds the full cropping strategies in Figure 5, highlighting the need for whole image analysis. Using SPE, IoU further improved, and ${{P}_{d}}$ reached 100%, which demonstrated the SPE is capable of fleshing out target shapes. The incorporation of the rough edge map raised the IOU further but led to target sticking. In M5, SPE is replaced with DPE. We can see that prompt embedding at the tail of the model highlights target edge differences and improves overall performance. Finally, BBM accurately eliminates false alarms and EGDSP achieves accurate annotation. Observe the visual maps in Figure 6. With the help of target energy initialization, our model quickly responds to the target area in the shallow layer and accurately predicts the left rotor of the drone. In scene two, without double prompt embedding (Ours w/o DPE), neighboring targets mix together. We further observe that the attention map also responded in the background region, revealing that the local shape prediction remains connected to the global context modeling. This finding underscores again the importance of preserving the entire scene. More details about each module are discussed in the supplementary material.
@@
We used the inferred pseudo masks from the test sets444 This dataset consists of 1079 samples, with the training and testing sets split 8:2. We performed three cross-validations. of SIRST, NUDT-SIRST, and IRSTD-1k to train the detection model, to validate the effectiveness of the pseudo labels for IRSTD. Quantitative results are shown in Table 4. EDGSP outperforms other single-point prompt label generation methods. Although MCGC achieves the lowest ${{F}_{a}}$, it also has a low IoU score. Remarkably, EDGSP not only reduces time cost by 80% but also outperforms full labels in detection task. Even with the coarse point prompt, EDGSP still achieves 99.5% performance of pixel-wise labeling. Visual results are shown in Figure 7. DNA-Net trained with EDGSP’s pseudo masks demonstrates enhanced shape and fewer false alarms compared to full labels. This reaffirms the applicability of the proposed methodology.
Figure 7: Visual results of DNA-Net. “Point Result” and “Full Result” indicate the outcomes of DNA-Net trained with the pseudo label from EDGSP and the full label, respectively. The correctly detected targets and false alarms are highlighted in orange and yellow circles, while incomplete shapes are marked with arrows.
@@ Conclusion
In this paper, we propose a novel annotation paradigm: enhancing the IRSTD network with the single-point prompt to accomplish high-precision IRSTLG. Specifically, we introduce an EDGSP framework. It includes target energy initialization for effective shape evolution, double prompt embedding to prevent label adhesion, and bounding box-based matching to eliminate false alarms. Three baselines equipped with EDGSP achieve accurate annotation on three IRSTD datasets for the first time. Excitingly, this is also the first time that a single point-generated pseudo mask outperforms full labels in practical IRSTD applications. In the future, it is worth investigating the coarse single-point prompt to further improve the utility of this framework.